# Monitoring Risks in Test-Time Adaptation

### v1

@@ 5 Experiments
  • $\hat{L}_{t}^{a}$: the estimated confidence lower bound on the running test risk under model adaptation (see Eq. 6). This direct extension of Podkopaev and Ramdas [30] preserves false-alarm guarantees but observes labels at each time point. While inapplicable in the unsupervised TTA setting, it serves as an oracle baseline. Since $L_{t}^{a}\geq L_{t}^{b}$ (under Assumption 1) our alarm $\Phi_{t}^{b}$ can never trigger before this alarm, $\Phi_{t}^{a}$, and consequently we inherit its detection delay. The closer $\hat{L}_{t}^{b}$ is to $\hat{L}_{t}^{a}$, the smaller is the price we pay for not observing test labels.
  • $\hat{L}_{t}^{c}$: a naive estimate of the running test risk, formed by substituting the supervised losses ${\mathbf{z}}_{0:t}$ with unsupervised proxies ${\mathbf{u}}_{0:t}$ in the alarm from Eq. 4 [30]. While it avoids using test labels, it lacks false alarm guarantees due to omitting the lower bounding step in Prop. 1.
  • $\hat{L}_{t}^{d}$: the estimated unsupervised lower bound on the running test risk of the static source model $p_{0}$ as presented in [1]. While providing false alarm guarantees without access to labels, it uses a different calibration procedure and is not applicable to a time-varying predictive model $p_{k}$.
@@
We monitor test risk using $0\text{-}1$ loss $\ell_{0\text{-}1}$ and Brier loss $\ell_{B}$ (Eq. 2). If not specified otherwise, we use a tolerance threshold of $\epsilon_{\text{tol}}=0.05$ for 00-$1$ loss and $\epsilon_{\text{tol}}=0.01$ for Brier loss. We set $\alpha=\alpha_{\text{source}}+\alpha_{\text{test}}$ to $0.2$ using most budget for controlling the test risk, i.e. $\alpha_{\text{test}}=0.175$ and $\alpha_{\text{source}}=0.025$. For threshold selection (§ 3.4) we use $N_{\text{cal}}=1000$ labeled samples from $P_{0}$.
@@
We evaluate our monitoring approach on three datasets: synthetic corruptions from ImageNet-C [12], and real-world distribution shifts from Yearbook [8] and FMoW-Time [45]. For ImageNet-C, we use the pretrained ViT-Base model [7] from the Timm library [34], focusing on Gaussian noise (GN) corruptions. Yearbook involves binary gender classification from portrait images, while FMoW-Time consists of satellite imagery with land use labels. Both datasets span multiple years; models are trained on data up to a cutoff year and tested on future samples. For Yearbook and FMoW, we follow the protocol of Yao et al. [45], using their provided model weights: a small CNN for Yearbook and DenseNet121 [14] for FMoW.
@@ 5.1 Illustrative Example
In the first experiment, we study the behavior of our alarm in comparison to described baselines. We are notably interested how closely our unsupervised monitoring tool mimics the two oracle quantities having access to the ground truth test labels: empirical running test risk $\hat{\bar{R}}_{t}$ and $\hat{L}_{t}^{a}$ [30]. To simulate an increasing test risk, we construct a test stream from ImageNet-C by gradually increasing the severity level of Gaussian noise corruption from in-distribution (no shift) up to severity level 5. We track both the unadapted source model and TENT using 00-$1$ loss $\ell_{0\text{-}1}$ and Brier loss $\ell_{B}$. We also verify the validity of Assumption 1 throughout adaptation by tracking $\Delta_{t}^{b}=PFP_{0}+\frac{1}{t}\sum_{k=1}^{t}PFN_{k}-PFP_{k}$. The assumption is met in practice when $\Delta_{t}^{b}\geq 0$ and violated when $\Delta_{t}^{b}<0$. $\Delta_{t}^{b}$ also reflects the tightness of $L_{t}^{b}$, so, ideally, it is also not too much above 00. We proceed analogously for Assumption 4.1 in Amoukou et al. [1] and denote it with $\Delta_{t}^{d}$.
The results are shown in Fig. 2. As the severity of the distribution shift increases, the empirical running test risk $\hat{\bar{R}}_{t}$ ( ) increases as well, for both 00-$1$ loss $\ell_{0\text{-}1}$ (top row) and Brier loss $\ell_{B}$ (bottom row). As expected, the unadapted source model (left) exhibits a higher risk, while adaptation with TENT (right) postpones the point where the empirical risk crosses the specified performance requirement. However, as the distribution shift becomes increasingly severe, $\hat{\bar{R}}_{t}$ eventually exceeds the upper bound on the source risk, $\hat{U}_{t}$, plus the tolerance margin $\epsilon_{\text{tol}}$ ( ), despite model adaptation. From this time point (  ), we wish to trigger an alarm. As expected, the lower confidence sequence on the empirical test risk, $\hat{L}_{t}^{a}$ ( ), which leverages test labels, detects the risk violation first. Encouragingly, our proposed unsupervised lower bound $\hat{L}_{t}^{b}$ ( ) closely follows the supervised bound $\hat{L}_{t}^{a}$. This indicates that our bound is tight and the price for not seeing labels is relatively small. The naive plugin bound, $\hat{L}_{t}^{c}$ ( ), is not only void of theoretical guarantees but also exhibits low power empirically by not detecting the risk violation in all but one case. The unsupervised bound by Amoukou et al. [1], $\hat{L}_{t}^{d}$ ( ), detects slightly later then $\hat{L}_{t}^{b}$ for Brier loss, but is extremely loose for $0\text{-}1$ loss. Fig. 2 shows that our Assumption 1 ( ) is met throughout the distribution shift in all cases, while the assumption of Amoukou et al. [1] ( ) is violated for $0\text{-}1$ loss, making their bound invalid for large $t$.
@@ 5.2 Generalization across Datasets, Shifts and TTA Methods
Next we evaluate the robustness of our monitoring tool by testing different TTA methods: TENT [42], CoTTA [41], SAR [27] and SHOT [19]. Please see § D.2 for details. We study four test streams: In-distribution of ImageNet (no shift, alarm should remain silent), ImageNet-C Gaussian noise (GN) severity level 5 (strong shift), Yearbook (moderate shift) and FMoW (gradual shift). Since classification error is the most commonly used metric in TTA, we track a risk increase for $0\textit{-}1$ loss.
@@
Fig. 9 shows that the empirical running test risk $\hat{\bar{R}}_{t}$ ( ) is closely mimicked by our $\hat{L}_{t}^{b}$ ( ) across TTA methods, datasets and shifts. Our alarm function correctly remains silent on the ID stream (first column) of ImageNet, where test risk remains below the threshold ( ). For FMoW, the risk increases steadily but also remains below the alarm threshold; this is accurately reflected in our monitoring, as $\hat{L}_{t}^{b}$ tightly tracks $\hat{\bar{R}}_{t}$ without triggering false alarms. For the immediate risk violation on ImageNet-C severity level 5 (second column), our test triggers an alarm after $<25$ steps for all TTA methods. Similar results are observed for Yearbook. SHOT results, consistent with Fig. 9, are shown in § A.2.
@@ 5.3 Detecting TTA Collapse
Unlike static models, the risk of a TTA model can increase not only due to distribution shift but also because the model deteriorates during adaptation. An extreme, yet well documented case of model failure in TTA is model collapse, where finally only a small subset or a single class is predicted [25, 39, 27, 22]. Alarmingly, these harmful collapses often occur silently [27]. We next ask whether our statistical framework can detect risk increases caused by model failure. This is not a given, as the monitor relies on the model’s own outputs (e.g., predictive uncertainty), which may become unreliable when the model itself fails. To induce model collapse, we follow [3] and apply TENT with a high learning rate of $\eta=1e^{-1}$ on the ImageNet-C (GN) corruption at severity level 1. We set a high $\epsilon_{\text{tol}}=0.2$ to disregard risk increase caused by distribution shift.
@@ 5.4 Alternative Loss Proxies
Figure 5: Comparison of loss proxies for last-layer TTA methods on ImageNet-C (GN) severity 5: Distance to class prototype is more effective than uncertainty for this TTA class.
In § 5.2, we demonstrated that using model uncertainty as a loss proxy yields valid (according to Assumption 1) and, importantly, tight unsupervised lower bounds across a representative set of TTA methods and data shifts. Here, we supplement these results with a case where relying solely on model uncertainty proves insufficient for effective detection. Specifically, when monitoring "last-layer" TTA methods [15, 35]—which adapt only the classification head $W=[{\bm{w}}_{1},\ldots,{\bm{w}}_{C}]\in\mathbb{R}^{H\times C}$—we observed that our unsupervised bound becomes overly loose, causing the alarm to fail under severe distribution shifts. We attribute this behavior to the normalization of each class prototype, i.e. ${\bm{w}}_{c}\>/\>||{\bm{w}}_{c}||_{2}^{2}$, at every adaptation step. This normalization leads to (much) reduced variability in uncertainty across samples at the start of adaptation, making the separation of high and low losses harder. To overcome this, we propose using the distance to the closest prototype [40, 23], $g({\mathbf{x}},p)=\min_{c}||f({\mathbf{x}})-{\bm{w}}_{c}||_{2}^{2}$, where $f$ denotes the feature extractor of model $p$, as an alternative loss proxy. Unlike uncertainty, this measure is less affected by the normalization of class prototypes. In Fig. 5, we show that this yields tighter lower bounds for T3A [15] and STAD [35]—both last-layer TTA methods—underscoring the importance of aligning the proxy choice with the specifics of the given TTA approach. Lastly, in § A.3 we show results when using energy score [21] as a loss proxy—finding that it yields looser bounds compared to model uncertainty.
@@ 6 Conclusion
While we have shown that our unsupervised alarm has detection delays not too much larger compared to its supervised counterpart [30] (see Fig. 2), the observed delays might still be too big for applications where detecting late is (significantly) more costly compared to raising false alarms. For such settings, it would be worth weakening the requirement on the probability of false alarm control under $H_{0}$ in order to gain more power under $H_{1}$. Perhaps this could be done by aiming for a weaker average run length control as is commonly done in the literature on change-point detection using confidence sequences [38, 37]. Moreover, although our proposed lower bound from Proposition 1 can be computed without access to test labels, verifying Assumption 1 for a given loss proxy still requires a labeled test stream. While we found empirically that this assumption holds in nearly all evaluated cases (Fig. 9), developing unsupervised diagnostics to flag potential violations of the assumption remains an important direction for future work.
@@
    • In § A.2, we extend results reported in § 5.2 with the TTA method SHOT.
@@ A.1 Improving Test Power
We also note that the scaled alarm function (Eq. 8) is closely related to the quantile detector proposed in Amoukou et al. [1]. Our work extends their approach in (at least) three main ways: first, by allowing for continuously evolving models, unlike [1] where a static model is assumed; second, by relaxing the assumption required for the loss proxy (see our Assumption 1 versus their Assumption 4.1); and third, by providing a theoretical justification for the increased reactivity of the high-probability test relative to the expected-loss test for continuous losses (via the cancellation of the loss threshold $\tau$). We elaborate further on these differences in § 4.
A.2 Results for SHOT
Figure 6: Estimated test risk for different datasets and TTA methods: We now also include results for SHOT to Fig. 9, which were omitted previously due to space constraints. The trend remains consistent, and risk violations are reliably detected for SHOT as well.
A.3 Alternative Loss Proxies
Here, we compare our choice of using model uncertainty as a loss proxy (§ 3.3) with the alternative of using the energy score [21], one of the most popular measures for detecting out-of-distribution (OOD) samples: $g({\mathbf{x}},p)=-\log\sum_{c=1}^{C}e^{m({\mathbf{x}})_{c}}$, where $m({\mathbf{x}})\in\mathbb{R}^{C}$ is a vector of logits for model $p$.
We present results in Fig. 7. While using energy as a loss proxy also yields valid lower bounds (as indicated by $\hat{\Delta}_{t}^{b}$ being positive), the resulting bounds are looser compared to those obtained using uncertainty as a proxy across all considered TTA methods and distribution shifts. We attribute the underperformance of the energy score to its primary focus on distinguishing between in-distribution and OOD inputs. In contrast, a useful loss proxy (see Assumption 1) should be able to differentiate between correctly predicted samples (i.e., low loss) and incorrectly predicted ones (i.e., high loss).
Although these findings further support our choice of model uncertainty as a loss proxy, we believe that exploring alternative proxies that would lead to (even) tighter bounds remains an important direction for future work.
@@ A.3 Alternative Loss Proxies
A.4 Ablation on Calibration Set Size
Our method requires a small labeled calibration set from the source distribution, which introduces additional computational overhead due to repeated evaluation of the adapted model on this set during adaptation. A small set of (labeled) source samples is commonly used for initializing TTA methods [52, 53, 41, 54, 55, 2, 56, 57, 58], and is indispensable for risk control [30, 1]. We next investigate whether the calibration set size can be reduced. In addition to addressing the reliance on labeled data, a smaller calibration set also reduces the runtime overhead of the online calibration procedure.

### v2

@@ 5 Experiments
  • $\hat{L}_{t}^{a}$: the estimated confidence lower bound on the running test risk under model adaptation (see Eq. 6). This direct extension of podkopaev2021tracking preserves false-alarm guarantees but observes labels at each time point. While inapplicable in the unsupervised TTA setting, it serves as an oracle baseline. Since $L_{t}^{a}\geq L_{t}^{b}$ (under Assumption 1) our alarm $\Phi_{t}^{b}$ can never trigger before this alarm, $\Phi_{t}^{a}$, and consequently we inherit its detection delay. The closer $\hat{L}_{t}^{b}$ is to $\hat{L}_{t}^{a}$, the smaller is the price we pay for not observing test labels.
  • $\hat{L}_{t}^{c}$: a naive estimate of the running test risk, formed by substituting the supervised losses ${\mathbf{z}}_{0:t}$ with unsupervised proxies ${\mathbf{u}}_{0:t}$ in the alarm from Eq. 4 [podkopaev2021tracking]. While it avoids using test labels, it lacks false alarm guarantees due to omitting the lower bounding step in Prop. 1.
  • $\hat{L}_{t}^{d}$: the estimated unsupervised lower bound on the running test risk of the static source model $p_{0}$ as presented in [jpmorgan2024sequential]. While providing false alarm guarantees without access to labels, it uses a different calibration procedure and is not applicable to a time-varying predictive model $p_{k}$.
@@
We monitor test risk using $0\text{-}1$ loss $\ell_{0\text{-}1}$ and Brier loss $\ell_{B}$ (Eq. 2). If not specified otherwise, we use a tolerance threshold of $\epsilon_{\text{tol}}=0.05$ for 0-11 loss and $\epsilon_{\text{tol}}=0.01$ for Brier loss. We set $\alpha=\alpha_{\text{source}}+\alpha_{\text{test}}$ to $0.2$ using most budget for controlling the test risk, i.e. $\alpha_{\text{test}}=0.175$ and $\alpha_{\text{source}}=0.025$. For threshold selection (§ 3.4) we use $N_{\text{cal}}=1000$ labeled samples from $P_{0}$.
@@
We evaluate our monitoring approach on three datasets: synthetic corruptions from ImageNet-C [hendrycks2019benchmarking], and real-world distribution shifts from Yearbook [ginosar2015century] and FMoW-Time [yao2022wild]. For ImageNet-C, we use the pretrained ViT-Base model [dosovitskiy2020image] from the Timm library [rw2019timm], focusing on Gaussian noise (GN) corruptions. Yearbook involves binary gender classification from portrait images, while FMoW-Time consists of satellite imagery with land use labels. Both datasets span multiple years; models are trained on data up to a cutoff year and tested on future samples. For Yearbook and FMoW, we follow the protocol of yao2022wild, using their provided model weights: a small CNN for Yearbook and DenseNet121 [huang2017densely] for FMoW.
@@ 5.1 Illustrative Example
In the first experiment, we study the behavior of our alarm in comparison to described baselines. We are notably interested how closely our unsupervised monitoring tool mimics the two oracle quantities having access to the ground truth test labels: empirical running test risk $\hat{\bar{R}}_{t}$ and $\hat{L}_{t}^{a}$ [podkopaev2021tracking]. To simulate an increasing test risk, we construct a test stream from ImageNet-C by gradually increasing the severity level of Gaussian noise corruption from in-distribution (no shift) up to severity level 5. We track both the unadapted source model and TENT using 0-11 loss $\ell_{0\text{-}1}$ and Brier loss $\ell_{B}$. We also verify the validity of Assumption 1 throughout adaptation by tracking $\Delta_{t}^{b}=PFP_{0}+\frac{1}{t}\sum_{k=1}^{t}PFN_{k}-PFP_{k}$. The assumption is met in practice when $\Delta_{t}^{b}\geq 0$ and violated when $\Delta_{t}^{b}<0$. $\Delta_{t}^{b}$ also reflects the tightness of $L_{t}^{b}$, so, ideally, it is also not too much above 0. We proceed analogously for Assumption 4.1 in jpmorgan2024sequential and denote it with $\Delta_{t}^{d}$.
The results are shown in Fig. 2. As the severity of the distribution shift increases, the empirical running test risk $\hat{\bar{R}}_{t}$ ( ) increases as well, for both 0-11 loss $\ell_{0\text{-}1}$ (top row) and Brier loss $\ell_{B}$ (bottom row). As expected, the unadapted source model (left) exhibits a higher risk, while adaptation with TENT (right) postpones the point where the empirical risk crosses the specified performance requirement. However, as the distribution shift becomes increasingly severe, $\hat{\bar{R}}_{t}$ eventually exceeds the upper bound on the source risk, $\hat{U}_{t}$, plus the tolerance margin $\epsilon_{\text{tol}}$ ( ), despite model adaptation. From this time point (  ), we wish to trigger an alarm. As expected, the lower confidence sequence on the empirical test risk, $\hat{L}_{t}^{a}$ ( ), which leverages test labels, detects the risk violation first. Encouragingly, our proposed unsupervised lower bound $\hat{L}_{t}^{b}$ ( ) closely follows the supervised bound $\hat{L}_{t}^{a}$. This indicates that our bound is tight and the price for not seeing labels is relatively small. The naive plugin bound, $\hat{L}_{t}^{c}$ ( ), is not only void of theoretical guarantees but also exhibits low power empirically by not detecting the risk violation in all but one case. The unsupervised bound by jpmorgan2024sequential, $\hat{L}_{t}^{d}$ ( ), detects slightly later then $\hat{L}_{t}^{b}$ for Brier loss, but is extremely loose for $0\text{-}1$ loss. Fig. 2 shows that our Assumption 1 ( ) is met throughout the distribution shift in all cases, while the assumption of jpmorgan2024sequential ( ) is violated for $0\text{-}1$ loss, making their bound invalid for large $t$.
@@ 5.2 Generalization across Datasets, Shifts and TTA Methods
Next we evaluate the robustness of our monitoring tool by testing different TTA methods: TENT [Wang2021TentFT], CoTTA [wang2022continual], SAR [niu2023towards] and SHOT [liang2020we]. Please see § D.2 for details. We study four test streams: In-distribution of ImageNet (no shift, alarm should remain silent), ImageNet-C Gaussian noise (GN) severity level 5 (strong shift), Yearbook (moderate shift) and FMoW (gradual shift). Since classification error is the most commonly used metric in TTA, we track a risk increase for $0\textit{-}1$ loss.
@@
Fig. 3 shows that the empirical running test risk $\hat{\bar{R}}_{t}$ ( ) is closely mimicked by our $\hat{L}_{t}^{b}$ ( ) across TTA methods, datasets and shifts. Our alarm function correctly remains silent on the ID stream (first column) of ImageNet, where test risk remains below the threshold ( ). For FMoW, the risk increases steadily but also remains below the alarm threshold; this is accurately reflected in our monitoring, as $\hat{L}_{t}^{b}$ tightly tracks $\hat{\bar{R}}_{t}$ without triggering false alarms. For the immediate risk violation on ImageNet-C severity level 5 (second column), our test triggers an alarm after <25<25 steps for all TTA methods. Similar results are observed for Yearbook. We additionally provide a detailed comparison with other baselines across all TTA methods in § A.4.
@@ 5.3 Detecting TTA Collapse
Unlike static models, the risk of a TTA model can increase not only due to distribution shift but also because the model deteriorates during adaptation. An extreme, yet well documented case of model failure in TTA is model collapse, where finally only a small subset or a single class is predicted [nguyen2023tipi, su2023beware, niu2023towards, marsden2024universal]. Alarmingly, these harmful collapses often occur silently [niu2023towards]. We next ask whether our statistical framework can detect risk increases caused by model failure. This is not a given, as the monitor relies on the model’s own outputs (e.g., predictive uncertainty), which may become unreliable when the model itself fails. To induce model collapse, we follow [boudiaf2022parameter] and apply TENT with a high learning rate of $\eta=1e^{-1}$ on the ImageNet-C (GN) corruption at severity level 1. We set a high $\epsilon_{\text{tol}}=0.2$ to disregard risk increase caused by distribution shift.
@@ 5.4 Alternative Loss Proxies
Figure 5: Comparison of loss proxies for last-layer TTA methods on ImageNet-C (GN) severity 5. Distance to class prototype is more effective than uncertainty for this TTA class.
[…]
@@ 6 Conclusion
While we have shown that our unsupervised alarm has detection delays not too much larger compared to its supervised counterpart [podkopaev2021tracking] (see Fig. 2), the observed delays might still be too big for applications where detecting late is (significantly) more costly compared to raising false alarms. For such settings, it would be worth weakening the requirement on the probability of false alarm control under $H_{0}$ in order to gain more power under $H_{1}$. Perhaps this could be done by aiming for a weaker average run length control as is commonly done in the literature on change-point detection using confidence sequences [shekhar2023sequential, shekhar2023reducing]. Moreover, although our proposed lower bound from Proposition 1 can be computed without access to test labels, verifying Assumption 1 for a given loss proxy still requires a labeled test stream. While we found empirically that this assumption holds in nearly all evaluated cases (Fig. 3), developing unsupervised diagnostics to flag potential violations of the assumption remains an important direction for future work.
@@ A.1 Improving Test Power
We also note that the scaled alarm function (Eq. 8) is closely related to the quantile detector proposed in jpmorgan2024sequential. Our work extends their approach in (at least) three main ways: first, by allowing for continuously evolving models, unlike [jpmorgan2024sequential] where a static model is assumed; second, by relaxing the assumption required for the loss proxy (see our Assumption 1 versus their Assumption 4.1); and third, by providing a theoretical justification for the increased reactivity of the high-probability test relative to the expected-loss test for continuous losses (via the cancellation of the loss threshold $\tau$). We elaborate further on these differences in § 4.
A.2 Alternative Loss Proxies
Here, we compare our choice of using model uncertainty as a loss proxy (§ 3.3) with two alternatives: energy score and entropy. The energy score [liu2020energy] is one of the most popular measures for detecting out-of-distribution (OOD) samples: $g({\mathbf{x}},p)=-\log\sum_{c=1}^{C}e^{m({\mathbf{x}})_{c}}$, where $m({\mathbf{x}})\in\mathbb{R}^{C}$ is a vector of logits for model $p$. The entropy of the predictive distribution is another common uncertainty-based measure: $g({\mathbf{x}},p)=-\sum_{c=1}^{C}p(c|{\mathbf{x}})\log p(c|{\mathbf{x}})$, where $p(c|{\mathbf{x}})$ is the predicted class probability. Entropy is widely used in TTA as an unsupervised objective, e.g., in TENT [Wang2021TentFT], where models adapt by minimizing predictive entropy on unlabeled test data.
We present the corresponding results in Fig. 6. While using energy as loss proxies also yields valid lower bounds (as indicated by $\hat{\Delta}_{t}^{b}>0$), the resulting bounds ( ) are consistently looser compared to those obtained when using model uncertainty ( ) across all TTA methods and distribution shifts. We attribute the underperformance of the energy score to its focus on distinguishing in-distribution versus OOD samples rather than separating correctly and incorrectly predicted instances. In contrast, a useful loss proxy (see Assumption 1) should be able to differentiate between correctly predicted samples (i.e., low loss) and incorrectly predicted ones (i.e., high loss).
For entropy ( ), detection delays are comparable to uncertainty as loss proxy. This confirms that our risk-monitoring mechanism remains valid and effective even when the same quantity is used both as the TTA objective and as the monitoring proxy. Indeed, our approach does not rely on the raw proxy values but instead on the proportion of test points exceeding a calibrated threshold (see Assumption 1). If adaptation minimizes entropy, our online calibration procedure (see § 3.4) dynamically adjusts this threshold, ensuring that the alarm remains reliable.
Although these findings further support our choice of model uncertainty as a suitable loss proxy, we believe that exploring alternative proxies that would lead to (even) tighter bounds remains an important direction for future work.
@@ A.2 Alternative Loss Proxies
A.3 Ablation on Calibration Set Size and Calibration Frequency
Our method requires a small labeled calibration set from the source distribution, which introduces additional computational overhead due to repeated evaluation of the adapted model on this set during adaptation. A small set of (labeled) source samples is commonly used for initializing TTA methods [song2023ecotta, lim2023ttn, wang2022continual, choi2022improving, adachi2023covariance, bar2024protected, jung2023cafa, lee2023tta, lee2024becotta], and is indispensable for risk control [podkopaev2021tracking, jpmorgan2024sequential]. We next investigate whether the calibration set size can be reduced. In addition to addressing the reliance on labeled data, a smaller calibration set also reduces the runtime overhead of the online calibration procedure.